A coin is tossed 3 times. Let Z is a number of tails. Find:
a) a probability density function f(z);
b) a cumulative distributional function F(z);
c) a mean E(Z) and a variance D(Z).
1
Expert's answer
2020-04-07T17:13:54-0400

(a)

We can consider all outcomes for this experiment: "\{ttt, thh, hth, hht 
, tth, tht, htt, hhh\}"

For "Z=0 : \ \{hhh\}"

"Z=1:\ \{thh, hth, hht\}"

"Z=2: \ \{ tth, tht,htt \}"

"Z=3: \ \{hhh\}"

Therefore, we have the probability distribution of "Z" :

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 z & 0 & 1 & 2 & 3 \\ \hline
 P(z) & \frac{1}{8} & \frac{3}{8} & \frac{3}{8} & \frac{1}{8}
\end{array}"

Probability density function:



(b)

"F(z)=P(Z\leq z)"

"F(0)=P(Z\leq 0)=P(0)=\frac{1}{8}, \ F(1)=P(Z\leq1)=P(0)+P(1)=\frac{1}{2}"

"F(2)=P(Z\leq 2)=P(0)+P(1)+P(2)=\frac{7}{8}, \ F(3)=P(Z\leq 3)=1"

"F(z)=\begin{cases}
\frac{1}{8}, \ z=0
\\ \frac{1}{2}, \ z=1
\\ \frac{7}{8}, \ z=2
\\ 1 , \ z=3
\end{cases}"

(c)

Mean value is "E(Z)=\sum zP(z)=0\times P(0)+1\times P(1)+2\times P(2)+3\times P(3)=0+\frac{3}{8}+\frac{6}{8} +\frac{3}8=\frac{12}{8}=1.5"


"D^2(Z)=[\sum z^2 P(z)]-E^2(Z)=[0^2\times P(0)+1^2\times P(1)+2^2\times P(2)+3^2\times P(3)-1.5^2=0+\frac{3}{8}+\frac{12}{8}+\frac{9}{8}-2.25=\frac{24}{8}-2.25=0.75"

Variance is "D(z)=\sqrt{0.75}\approx 0.866"